OpenStudy (anonymous):

Solve for the real value of x: 5/e^x+1=1

OpenStudy (jdoe0001):

hmmm have you covered logarithms yet?

OpenStudy (anonymous):

yes but it was so long ago i dont remember . i know it can be converted to natural log

OpenStudy (joannablackwelder):

Is it \[\frac{ 5 }{ e ^{x} +1}=1\]

OpenStudy (anonymous):

yes !

OpenStudy (joannablackwelder):

Ok. First get x in the numerator by multiplying by the denominator on both sides.

OpenStudy (joannablackwelder):

That leaves us with. \[5=e ^{x}+1\]

OpenStudy (joannablackwelder):

Following me so far?

OpenStudy (anonymous):

did you cross multiply ?

OpenStudy (joannablackwelder):

That is one way to describe it. :)

OpenStudy (anonymous):

okay i understand then , continue ?

OpenStudy (joannablackwelder):

Good. Now, continue working on isolating x by subtracting 1 on both sides.

OpenStudy (joannablackwelder):

Can you tell me what that leaves us with?

OpenStudy (anonymous):

[5-1=e^{x}\]

OpenStudy (joannablackwelder):

Right or \[4=e ^{x}\]

OpenStudy (anonymous):

yeah sorry about the brackets so do we have to convert to natural log ?

OpenStudy (joannablackwelder):

No worries. Yes, take the natural log of both sides.

OpenStudy (joannablackwelder):

What did you get?

OpenStudy (anonymous):

does e = ln ?

OpenStudy (joannablackwelder):

No, but ln(e^x)=x Natural log cancels out e^

OpenStudy (anonymous):

is there a rule to use any time you're converting from e to ln or ln to e ?

OpenStudy (joannablackwelder):

I'm using the idea that ln(e)=1

OpenStudy (anonymous):

im confused on how you convert then . how can you go from e^x to ln if when you times them they equal 1 ?

OpenStudy (anonymous):

i understand you up to this point :(

OpenStudy (joannablackwelder):

Just a sec.

OpenStudy (xapproachesinfinity):

well one way to think about it is that lnx and e^x are inverse of each other fo f(g(x))=x comes in handy here whether you do this ln(e^x)=x or e^(lnx)=x you always end up with the identity function x

OpenStudy (joannablackwelder):

My bad. I gave you the wrong property. It should be \[\log _{b}b ^{x}=x\]

OpenStudy (joannablackwelder):

So that brings us to ln4=x :)

OpenStudy (xapproachesinfinity):

lne=1 since e^1=e. you always thing of log and exponential as inverses in other words they undo each other

OpenStudy (xapproachesinfinity):

you can also rewrite to make it more funcy 2ln2=x do you know what property is this?

OpenStudy (anonymous):

why are there two b's in joannas formula ? and no i dont

OpenStudy (xapproachesinfinity):

that logarithms base b ln is logar base e

OpenStudy (joannablackwelder):

The base and what you are taking the log of are the same.

OpenStudy (joannablackwelder):

Natural log is log base e.

OpenStudy (xapproachesinfinity):

 now think of this power that has base 2

OpenStudy (xapproachesinfinity):

we went from 2 to the power of 3=8 log helps us to undo this powering if u may say

OpenStudy (xapproachesinfinity):